# Automated Layout and Control Co-Design of Robust Multi-UAV Transportation Systems

### v3

@@ II Methodology
The co-design problem we solve is intrinsically twofold, as it involves optimizing over a physical configuration as well as over control parameters. The question we ask is: what is the physical placement of $N$ quadcopters around a payload, and associated controller, such that, the overall flying system is the most robust possible to disturbances? We cast the problem as an optimization problem where the cost function is inspired by $\mathcal{H}_{2}$ control theory, and the decision variables are the quadcopters’ attachment locations. In the inner loop of the optimization, a Linear Quadratic Regulator (or LQR, i.e. the $\mathcal{H}_{2}$-optimal controller) is associated to candidate layouts for evaluation. This choice, in fact, allows a straightforward computation of the control (i.e. a linear feedback matrix), and also an analytic evaluation of the $\mathcal{H}_{2}$-inspired cost function. More advanced control strategies could be used, at the expense of a more involved policy optimization, and a more computationally expensive evaluation of the cost function (using for example Monte Carlo methods). We make the assumption that a reliable system for state estimation is available.
@@ II-A Modelling
Each quadcopter is seen as a thrust module providing four inputs to the system. Therefore, if $N$ quadcopters are deployed, the thrust input vector is $4N$-dimensional, i.e. $\mathbf{u}\in\mathbb{R}^{4N}$. We assume that each thrust component is limited to the interval $[u_{l},u_{h}]\subset\mathbb{R}$.
@@
Figure 2: a) Extrusion payload geometry and representation of the mid-plane. Its intersection with the side faces is the curve $\Gamma$. b) Schematics of four quadcopters attached around the payload along $\Gamma$. Representation of the local body reference frame centered at the centroid of $\Gamma$, placement variables defined by the angles $\boldsymbol{\theta_{i}}$ ($i=1,...,4$), and disturbance force $\mathbf{f}_{d}$ and torque $\mathbf{t}_{d}$.
@@
The dynamics matrix $A\in\mathbb{R}^{12\times 12}$ is obtained by taking the Jacobian of (1) around hovering conditions with respect to the state vector $\mathbf{x}$. $A$ does not depend on the placement variables $\boldsymbol{\theta}$. The matrices $B(\boldsymbol{\theta})\in\mathbb{R}^{12\times 4N}$ and $B_{d}(\boldsymbol{\theta})\in\mathbb{R}^{12\times 6}$ are obtained by taking the Jacobian of (1) with respect to the input $\mathbf{u}$ and disturbance $\mathbf{d}$ vectors, respectively. These matrices depend on $\boldsymbol{\theta}$ directly or through the inertia of the system (which itself is a function of $\boldsymbol{\theta}$). It is also important to note that the error due to the linearization of the dynamics equations does not depend on $\boldsymbol{\theta}$, and the linear approximation is reasonably accurate for the applications of interest in this paper.
@@ II-B Control and layout co-design
[…]
@@
$d_{M}(\mathbf{u};\bar{\mathbf{u}},\Sigma_{\mathbf{u}})=\sqrt{(\mathbf{u}-\bar{\mathbf{u}})\Sigma_{\mathbf{u}}^{-1}(\mathbf{u}-\bar{\mathbf{u}})}.$ (10)
@@
[…]
@@
weight $\text{weight matrices}\,\,C\,\,\text{and}\,\,D$
@@
     Compute $A$, $B(\boldsymbol{\theta})$, $B_{d}(\boldsymbol{\theta})$ and feedforward $\bar{\mathbf{u}}$
@@ II-C Control Implementation
For simplicity, we treat the integrated quadcopter-payload system as a single rigid body, and use the LQR gains computed in the co-design optimization algorithm 1, which are readily available as a byproduct of our co-design tool. This approach leads to a hierarchical control infrastructure, in which a single estimator is used for the overall motion and the quadcopters receive thrust commands individually through $N$ parallel communication channels. A block diagram of the control infrastructure is shown in Fig. 4.
@@
The overall behavior of the controller is dominated by the choice of some hyperparameters which directly impact the computation of the $\mathcal{H}_{2}$ cost. First, the disturbance matrix $B_{d}$ from (2), that describes how the random disturbance $\mathbf{d}$ affects the dynamics, and, in particular, quantifies its covariance along different directions in the wrench space. Second, the matrices $C$ and $D$ from (3), i.e. how the auxiliary variable $z$ is defined. For simplicity, the $B_{d}$ matrix was chosen to weigh the six disturbance components equally, i.e. $B_{d}(4,1)=B_{d}(5,2)=B_{d}(6,3)=B_{d}(10,4)=B_{d}(11,5)=B_{d}(12,6)=1$. The $C$ matrix is chosen to just weigh the position of the payload centroid and the system’s yaw angle (i.e. $C\in\mathbb{R}^{(4+4N)\times 12}$). After evaluating the system’s behavior in simulation, we set $C(1,1)=C(2,2)=0.5$ (corresponding to the position on a horizontal plane $x-y$), $C(3,3)=10$ (corresponding to vertical position $z$), $C(4,9)=50$ (corresponding to yaw angle), and zero for all the other components. The input weight matrix $D\in\mathbb{R}^{(4+4N)\times 4N}$ was set to the block matrix $D=[\mathbf{0}_{4N\times 4},I_{4N}]^{T}$, where $\mathbf{0}_{m\times n}$ denotes the $m\times n$ matrix of all zero elements and $I_{k}$ the $k$-dimensional identity.
@@ II-D Optimization Results
quadcopter size (motor to motor) $330\,\mathrm{mm}$ propeller diameter $114.5\,\mathrm{mm}$
@@ III Experiments
The relevant parameters of the vehicles used are shown in Tab. I. The battery capacity ($5300\,\mathrm{mAh}$) has been chosen to guarantee $10$ minutes of flight time at maximum payload. The system theoretically is able to lift an external payload of up to $\sim 0.8\,\mathrm{kg}$ per quadcopter. It is interesting to note that the main contribution to the overall inertia of the system is due to the quadcopters, since they are attached to the payload perimeter, and thus at a larger distance from the overall center of mass. It is also important to highlight that contributing a greater inertia than the actual payload is in line with all the ground and aerial transportation systems in use nowadays. In fact, most of the transportation systems that we use daily (e.g. trains, planes, cars) have significantly greater mass and inertia compared to the payload they transport.
@@ III-A Hovering
As a first comparison, we report the hovering performances of the different systems in the various configurations we tested. The hovering was performed in still air and with a $1\,\mathrm{m\cdot s^{-1}}$ wind disturbance. The RMSE values, averaged across five runs roughly $15$ seconds long, of position deviation from reference and attitude angles can be found in Tabs. II and III. The optimal configurations for both panels are more stable and less subject to disturbance. The difference is particularly significant for Panel (B), which could not withstand the wind disturbance at all, and showed an unstable behavior (due to actuator saturation) in all the five trials.
@@ III-D Disturbance Rejection
Figure 7: $x-z$ positions of the optimal and suboptimal configurations on Panel (A). The position reference step consists of an offset of $1\,\mathrm{m}$ along $x$. The optimal configuration has lower overshoot and faster recovery.

### v4

@@ II Methodology
The co-design problem we solve is intrinsically twofold, as it involves optimizing over a physical configuration as well as over control parameters. The question we ask is: what is the physical placement of NN quadcopters around a payload, and associated controller, such that, the overall flying system is the most robust possible to disturbances? We cast the problem as an optimization problem where the cost function is inspired by $\mathcal{H}_{2}$ control theory, and the decision variables are the quadcopters’ attachment locations. In the inner loop of the optimization, a Linear Quadratic Regulator (or LQR, i.e. the $\mathcal{H}_{2}$-optimal controller) is associated to candidate layouts for evaluation. This choice, in fact, allows a straightforward computation of the control (i.e. a linear feedback matrix), and also an analytic evaluation of the $\mathcal{H}_{2}$-inspired cost function. More advanced control strategies could be used, at the expense of a more involved policy optimization, and a more computationally expensive evaluation of the cost function (using for example Monte Carlo methods). We make the assumption that a reliable system for state estimation is available.
@@ II-A Modelling
Each quadcopter is seen as a thrust module providing four inputs to the system. Therefore, if NN quadcopters are deployed, the thrust input vector is $4N$-dimensional, i.e. $\mathbf{u}\in\mathbb{R}^{4N}$. We assume that each thrust component is limited to the interval $u_{l},u_{h}]\subset\mathbb{R}$.
@@
Figure 2: a) Extrusion payload geometry and representation of the mid-plane. Its intersection with the side faces is the curve $\Gamma$. b) Schematics of four quadcopters attached around the payload along $\Gamma$. Representation of the local body reference frame centered at the centroid of $\Gamma$, placement variables defined by the angles $\boldsymbol{\theta_{i}}$ (i=1,,4i=1,...,4), and disturbance force $\mathbf{f}_{d}$ and torque $\mathbf{t}_{d}$.
@@
The dynamics matrix $A\in\mathbb{R}^{12\times 12}$ is obtained by taking the Jacobian of (1) around hovering conditions with respect to the state vector $\mathbf{x}$. AA does not depend on the placement variables $\boldsymbol{\theta}$. The matrices $B(\boldsymbol{\theta})\in\mathbb{R}^{12\times 4N}$ and $B_{d}(\boldsymbol{\theta})\in\mathbb{R}^{12\times 6}$ are obtained by taking the Jacobian of (1) with respect to the input $\mathbf{u}$ and disturbance $\mathbf{d}$ vectors, respectively. These matrices depend on $\boldsymbol{\theta}$ directly or through the inertia of the system (which itself is a function of $\boldsymbol{\theta}$). It is also important to note that the error due to the linearization of the dynamics equations does not depend on $\boldsymbol{\theta}$, and the linear approximation is reasonably accurate for the applications of interest in this paper.
@@ II-B Control and layout co-design
and minimize the integral $\mathcal{H}_{2}$ control cost $J=\int_{0}^{\infty}\mathbf{z}^{T}\mathbf{z}\,dt$ when the system is subject to the Gaussian disturbance $\mathbf{d}$. In other words, the goal here is to capture the first-order effect of disturbances such as random forces and torques due to a wind gust. The linear approximation allows us to capture this first-order effect in a computationally tractable way, thanks to the following analytic derivation. This derivation would be less suitable for higher-order nonlinear metrics aiming at capturing, for example, the system maneuverability or maximum achievable speed.
@@
$d_{M}(\mathbf{u};\bar{\mathbf{u}},\Sigma_{\mathbf{u}})=\sqrt{(\mathbf{u}-\bar{\mathbf{u}})^{T}\Sigma_{\mathbf{u}}^{-1}(\mathbf{u}-\bar{\mathbf{u}})}.$ (10)
@@
$\text{s.t.}\,\,\mathbf{u}_{l}\leq\mathbf{u}\leq\mathbf{u}_{h}\,\,\,\text{(element-wise)}$
@@
$\text{weight matrices}\,\,C\,\,\text{and}\,\,D$
@@
  Compute AA, $B(\boldsymbol{\theta})$, $B_{d}(\boldsymbol{\theta})$ and feedforward $\bar{\mathbf{u}}$
@@ II-C Control Implementation
For simplicity, we treat the integrated quadcopter-payload system as a single rigid body, and use the LQR gains computed in the co-design optimization algorithm 1, which are readily available as a byproduct of our co-design tool. This approach leads to a hierarchical control infrastructure, in which a single estimator is used for the overall motion and the quadcopters receive thrust commands individually through NN parallel communication channels. A block diagram of the control infrastructure is shown in Fig. 4.
@@
The overall behavior of the controller is dominated by the choice of some hyperparameters which directly impact the computation of the $\mathcal{H}_{2}$ cost. First, the disturbance matrix $B_{d}$ from (2), that describes how the random disturbance $\mathbf{d}$ affects the dynamics, and, in particular, quantifies its covariance along different directions in the wrench space. Second, the matrices $C$ and $D$ from (3), i.e. how the auxiliary variable $z$ is defined. For simplicity, the $B_{d}$ matrix was chosen to weigh the six disturbance components equally, i.e. $B_{d}(4,1)=B_{d}(5,2)=B_{d}(6,3)=B_{d}(10,4)=B_{d}(11,5)=B_{d}(12,6)=1$. The $C$ matrix is chosen to just weigh the position of the payload centroid and the system’s yaw angle (i.e. $C\in\mathbb{R}^{(4+4N)\times 12}$). After evaluating the system’s behavior in simulation, we set $C(1,1)=C(2,2)=0.5$ (corresponding to the position on a horizontal plane xyx-y), $C(3,3)=10$ (corresponding to vertical position $z$), $C(4,9)=50$ (corresponding to yaw angle), and zero for all the other components. The input weight matrix $D\in\mathbb{R}^{(4+4N)\times 4N}$ was set to the block matrix $D=[\mathbf{0}_{4N\times 4},I_{4N}]^{T}$, where $\mathbf{0}_{m\times n}$ denotes the $m\times n$ matrix of all zero elements and $I_{k}$ the $k$-dimensional identity.
@@ II-D Optimization Results
quadcopter size (motor to motor) $330\,\mathrm{mm}$ propeller diameter $203\,\mathrm{mm}$
@@ III Experiments
The relevant parameters of the vehicles used are shown in Tab. I. The battery capacity ($5300\,\mathrm{mAh}$) has been chosen to guarantee 1010 minutes of flight time at maximum payload. The system theoretically is able to lift an external payload of up to $\sim 0.8\,\mathrm{kg}$ per quadcopter. It is interesting to note that the main contribution to the overall inertia of the system is due to the quadcopters, since they are attached to the payload perimeter, and thus at a larger distance from the overall center of mass. It is also important to highlight that contributing a greater inertia than the actual payload is in line with all the ground and aerial transportation systems in use nowadays. In fact, most of the transportation systems that we use daily (e.g. trains, planes, cars) have significantly greater mass and inertia compared to the payload they transport.
@@ III-A Hovering
As a first comparison, we report the hovering performances of the different systems in the various configurations we tested. The hovering was performed in still air and with a $1\,\mathrm{m\cdot s^{-1}}$ wind disturbance. The RMSE values, averaged across five runs roughly 1515 seconds long, of position deviation from reference and attitude angles can be found in Tabs. II and III. The optimal configurations for both panels are more stable and less subject to disturbance. The difference is particularly significant for Panel (B), which could not withstand the wind disturbance at all, and showed an unstable behavior (due to actuator saturation) in all the five trials.
@@ III-D Disturbance Rejection
Figure 7: xzx-z positions of the optimal and suboptimal configurations on Panel (A). The position reference step consists of an offset of $1\,\mathrm{m}$ along $x$. The optimal configuration has lower overshoot and faster recovery.